OpenStudy (anonymous):

A machine grinds 2/7 pounds of spices in 1/3 second. What is the machines rate in pounds per second of grinding spices

OpenStudy (anonymous):

Well, if you know it grinds 2/7lbs of spices in 1/3 of a second, you need to find how many lbs it can grind in 1 second. You can do this by multiplying both sides by 3: \[\frac{ 2 }{ 7 } = \frac{ 1 }{ 3 }\] \[\frac{ 2 }{ 7 } \times 3 = \frac{ 6 }{ 7 }\] \[\frac{ 1 }{ 3 } \times 3 = 1\] Therefore, \[\frac{ 6 }{ 7 } = 1\] The machine grinds 6/7 lbs of spices per second. Hope this helped!